# Second-Order Convergence in Private Stochastic Non-Convex Optimization

### v1

@@ Abstract
We investigate the problem of finding second-order stationary points (SOSP) in differentially private (DP) stochastic non-convex optimization. Existing methods suffer from two key limitations: (i) inaccurate convergence error rate due to overlooking gradient variance in the saddle point escape analysis, and (ii) dependence on auxiliary private model selection procedures for identifying DP-SOSP, which can significantly impair utility, particularly in distributed settings. To address these issues, we propose a generic perturbed stochastic gradient descent (PSGD) framework built upon Gaussian noise injection and general gradient oracles. A core innovation of our framework is using model drift distance to determine whether PSGD escapes saddle points, ensuring convergence to approximate local minima without relying on second-order information or additional DP-SOSP identification. By leveraging the adaptive DP-SPIDER estimator as a specific gradient oracle, we develop a new DP algorithm that rectifies the convergence error rates reported in prior work. We further extend this algorithm to distributed learning with arbitrarily heterogeneous data, providing the first formal guarantees for finding DP-SOSP in such settings. Our analysis also highlights the detrimental impacts of private selection procedures in distributed learning under high-dimensional models, underscoring the practical benefits of our design. Numerical experiments on real-world datasets validate the efficacy of our approach.
@@ 1 Introduction
3. Application to distributed learning: We extend the adaptive DP-SPIDER estimator to distributed learning. Via adaptivity, our learning algorithm improves upon the DIFF2 [36], which only guarantees convergence to DP-FOSP under homogeneous data. In contrast, our method provides the first error bound for converging to DP-SOSP under arbitrarily heterogeneous data: $\tilde{O}\big{(}\frac{1}{(mn)^{1/3}}+\big{(}\frac{\sqrt{d}}{\epsilon mn}\big{)}^{2/5}\big{)}$, where $m$ is the number of participants and $n$ is the number of samples per participant. Furthermore, we analyze the adverse effects of private model selection, showing that it deteriorates utility guarantees in high-dimensional regimes, thereby highlighting the necessity of our framework.
@@ 2 Related Work
With the rise of large-scale models and decentralized data, distributed learning has gained significant attention. Methods like federated learning [33] have enabled multiple clients to collaboratively train models without sharing their local data. Recent studies, such as [15, 52, 31, 32] have investigated DP learning in distributed settings, but these works are limited to first-order optimality. While some studies have investigated SOSP in distributed learning, their focus was primarily on Byzantine-fault tolerance [54], and communication efficiency [35, 6]. No effort, to our knowledge, has been made to ensure DP-SOSP in distributed learning scenarios with heterogeneous data. Our proposed framework fills this gap by introducing the first distributed learning algorithm with DP-SOSP guarantees while effectively handling arbitrary data heterogeneity across clients.
@@ 4 Preliminaries
Distributed Learning. In the distributed (federated) learning setting, $m$ clients collaboratively learn under the coordination of a central server. Each client $j\in[m]$ has a local dataset $D_{j}$ of size $n$, sampled from an unknown local distribution $\mathcal{D}_{j}$. The population risk for client $j$ is defined as $F_{\mathcal{D}_{j}}(x)\coloneqq\mathbb{E}_{z\sim\mathcal{D}j}[f(x;z)]$ or simply $F_{j}(x)$. The global population risk is defined as the average of the local population risks: $F_{\mathcal{D}}(x)\coloneqq\frac{1}{m}\sum_{j\in[m]}F_{j}(x)$, or simply $F(x)$. We allow for heterogeneous local datasets, meaning that the local distributions $\{\mathcal{D}_{j}\}_{j\in[m]}$ may differ arbitrarily.
@@ 5.2 Our Approach: A General Gaussian-Perturbed SGD Framework
[…]
@@ 5.3 Main Results for Gauss-PSGD Framework
Here $C$ is an absolute constant that may change across expressions. Let $\tilde{x}$ denote a saddle point of the population risk $F(\cdot)$, and $\mathcal{H}\coloneqq\nabla^{2}F(\tilde{x})$. Let $v_{\min}$ be the eigenvector corresponding to $\lambda_{\min}(\mathcal{H})$, and $\mathcal{P}_{-v_{\min}}$ be the projection onto the orthogonal complement of $v_{\min}$. Set $\gamma\coloneqq-\lambda_{\min}(\mathcal{H})$.
@@
With probability at least $21-\omega/$ 2, for all $t\in[T]$, $\|\nu_{t}\|\leq C\sqrt{2\log\left(\frac{4T}{\omega}\right)}\psi\leq\chi$.
@@ 6.2 Error Rate Analysis for DP-SOSP with Ada-DP-SPIDER
             /* Using oracle $\mathcal{O}_{1}$/
@@ 9             $\textsf{drift}\leftarrow 0$;
             /* Using oracle $\mathcal{O}_{2}$/
@@ 7 Extension to Distributed SGD
The learning algorithm using Distributed Ada-DP-SPIDER can be viewed as an adaptive extension of the DIFF2 algorithm [36], which uses standard SPIDER and is limited to convergence to DP-FOSP under homogeneous data. To the best of our knowledge, our method is the first to achieve convergence to a DP-SOSP in a distributed setting with arbitrarily heterogeneous data.
@@ A.2 Privacy Preliminaries
Given any input data $D\in\mathcal{X}^{n}$ and a query function $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, the Gaussian mechanism $\mathcal{M}_{G}$ is defined as $q(D)+\nu$ where $\nu\sim\mathcal{N}(0,\sigma_{G}^{2}\mathbf{I}_{d})$. Let $\Delta_{2}(q)$ be the $\ell_{2}$-sensitivity of $q$, i.e., $\Delta_{2}(q)\coloneqq\sup_{D\sim D^{\prime}}\|q(D)-q(D^{\prime})\|_{2}$. For any $\sigma,\delta>0$, $\mathcal{M}_{G}$ guarantees $(\frac{\Delta_{2}(q)}{\sigma_{G}}\sqrt{2\log\frac{1.25}{\delta}},\delta)$-DP. That is, if we want the output of $q$ to be $(\epsilon,\delta)$-DP for any $10<\epsilon,\delta<$ 1, then $\sigma_{G}$ should be set to $\frac{\Delta_{2}(q)}{\epsilon}\sqrt{2\log\frac{1.25}{\delta}}$.
@@
Given target privacy parameters $10<\epsilon<$ 1 and $10<\delta<$ 1, to ensure $(\epsilon,\delta)$-DP over $k$-fold adaptive mechanisms, it suffices that each mechanism is $(\epsilon^{\prime},\delta^{\prime})$-DP, where $\epsilon^{\prime}=\frac{\epsilon}{2\sqrt{2k\ln(2/\delta)}}$ and $\delta^{\prime}=\frac{\delta}{2k}$.
@@ B.6 Proof of Lemma 5
Applying a union bound over $t\in[T]$ gives the desired result: with probability at least $21-\omega/$ 2, $\|\hat{g}_{t}-\nabla F(x_{t-1})\|\leq\chi$ for all $t$. ∎
@@ B.7 Proof of Lemma 6
By Lemma 5, with probability at least $21-\omega/$ 2, the gradient estimation error satisfies $\|\hat{g}_{t}-\nabla F(x_{t-1})\|\leq\chi$ for all $t\in[T]$. We analyze two cases based on whether the algorithm is in the escape phase.
@@
By Corollary 1, each $\Gamma$-descent succeeds with probability at least $1/8$, and we boost this to $21-\omega/$ 2 via the $Q$ independent repetitions in every escape procedure. By Lemma 2 with failure probability $\omega_{0}=\frac{\omega}{2N_{\text{saddle}}}$, we require:
@@ C.2 Proof of Lemma 8
By Lemma 5, with probability at least $21-\omega/$ 2, we have $\|\nabla F(x_{t-1})-\hat{g}_{t}\|\leq\chi$ for all $t$.
@@
$-U\leq-\frac{|\mathcal{T}|\kappa}{6\eta}+5T\eta\chi^{2},$ (97)
@@ D.3 Proof of Theorem 4
Gradient Estimation Error. For any $\mathcal{S}_{j}$ and $x$, $\nabla\hat{f}_{\mathcal{S}_{j}}(x)-\nabla F_{j}(x)$ is zero-mean and follows $\operatorname{nSG}\left(\frac{2G}{\sqrt{n}}\right)$. By the $G$-Lipschitz assumption and norm-sub-Gaussian concentration (Lemma 11), we have with probability at least $81-\omega^{\prime}/$ 8:
@@
Also, since $\theta_{p}\sim\mathcal{N}(0,\sigma_{1}^{2}/m)$, standard Gaussian concentration (Lemma 10) gives, with probability at least $81-\omega^{\prime}/$ 8:
@@
Hessian Estimation Error. For any $j\in[m]$ and $z\in\mathcal{S}_{j}$, $\mathbb{E}[\nabla^{2}f(x_{p};z)-\nabla^{2}F_{j}(x_{p})]=0$, and $\|\nabla^{2}f(x_{p};z)-\nabla^{2}F_{j}(x_{p})\|_{2}\leq 2M$ (due to $M$-smoothness). That is, each empirical Hessian term is $2M$-bounded in operator norm. Applying the matrix Bernstein inequality (Lemma 14), and using the assumption $mn\geq\frac{4}{9}\log(8d/\omega^{\prime})$, we obtain with probability at least $81-\omega^{\prime}/$ 8:
@@
For the added noise, since $\mathbf{H}_{p}$ consists of symmetric Gaussian matrices with variance $\sigma_{2}^{2}/m$, Lemma 15 gives, with probability at least $81-\omega^{\prime}/$ 8:
@@
Verification for $x_{p}$. Combining the above estimates and using a union bound, with probability at least $21-\omega^{\prime}/$ 2, we have:
@@
[…]
@@
Hence, $x_{p}$ will be selected with probability at least $21-\omega^{\prime}/$ 2.

### v2

@@ Abstract
We investigate the problem of finding second-order stationary points (SOSP) in differentially private (DP) stochastic non-convex optimization. Existing methods suffer from two key limitations: (i) inaccurate convergence error rate due to overlooking gradient variance in the saddle point escape analysis, and (ii) dependence on auxiliary private model selection procedures for identifying DP-SOSP, which can significantly impair utility, particularly in distributed settings. To address these issues, we propose a generic perturbed stochastic gradient descent (PSGD) framework built upon Gaussian noise injection and general gradient oracles. A core innovation of our framework is using model drift distance to determine whether PSGD escapes saddle points, ensuring convergence to approximate local minima without relying on second-order information or additional DP-SOSP identification. By leveraging the adaptive DP-SPIDER estimator as a specific gradient oracle, we develop a new DP algorithm that rectifies the convergence error rates reported in prior work. We further extend this algorithm to distributed learning with heterogeneous data, providing the first formal guarantees for finding DP-SOSP in such settings. Our analysis also highlights the detrimental impacts of private selection procedures in distributed learning under high-dimensional models, underscoring the practical benefits of our design. Numerical experiments on real-world datasets validate the efficacy of our approach.
@@ 1 Introduction
3. Application to distributed learning: We extend the adaptive DP-SPIDER estimator to distributed learning. Via adaptivity, our learning algorithm improves upon the DIFF2 [36], which only guarantees convergence to DP-FOSP under homogeneous data. In contrast, our method provides the first error bound for converging to DP-SOSP under heterogeneous data: $\tilde{O}\big(\frac{1}{(mn)^{1/3}}+\big(\frac{\sqrt{d}}{\epsilon mn}\big)^{2/5}\big)$, where $m$ is the number of participants and $n$ is the number of samples per participant. Furthermore, we analyze the adverse effects of private model selection, showing that it deteriorates utility guarantees in high-dimensional regimes, thereby highlighting the necessity of our framework.
@@ 2 Related Work
With the rise of large-scale models and decentralized data, distributed learning has gained significant attention. Methods like federated learning [33] have enabled multiple clients to collaboratively train models without sharing their local data. Recent studies, such as [15, 52, 31, 32] have investigated DP learning in distributed settings, but these works are limited to first-order optimality. While some studies have investigated SOSP in distributed learning, their focus was primarily on Byzantine-fault tolerance [54], and communication efficiency [35, 6]. No effort, to our knowledge, has been made to ensure DP-SOSP in distributed learning scenarios with heterogeneous data. Our proposed framework fills this gap by introducing the first distributed learning algorithm with DP-SOSP guarantees while effectively handling data heterogeneity across clients.
@@ 4 Preliminaries
Distributed Learning. In the distributed (federated) learning setting, $m$ clients collaboratively learn under the coordination of a central server. Each client $j\in[m]$ has a local dataset $D_{j}$ of size $n$, sampled from an unknown local distribution $\mathcal{D}_{j}$. The population risk for client $j$ is defined as $F_{\mathcal{D}_{j}}(x)\coloneqq\mathbb{E}_{z\sim\mathcal{D}j}[f(x;z)]$ or simply $F_{j}(x)$. The global population risk is defined as the average of the local population risks: $F_{\mathcal{D}}(x)\coloneqq\frac{1}{m}\sum_{j\in[m]}F_{j}(x)$, or simply $F(x)$. We allow for heterogeneous local datasets, meaning that the local distributions $\{\mathcal{D}_{j}\}_{j\in[m]}$ may differ.
@@ 5.2 Our Approach: A General Gaussian-Perturbed SGD Framework
We present our Gauss-PSGD framework in Algorithm 1. As specified in (2), we employ a general Gaussian-perturbed stochastic gradient oracle, denoted as $\texttt{P\_Grad\_Oracle}(*)$ in steps 4 and 10, where * abstracts the specific arguments required by the oracle implementation. This abstraction allows Gauss-PSGD to serve as a flexible optimization framework for non-convex stochastic problems, applicable beyond the differential privacy (DP) setting.
At each iteration, the gradient estimate $\hat{g}_{t}$ is computed by $\texttt{P\_Grad\_Oracle}(*)$, and the model parameter is updated via the gradient descent step in (4). The algorithm proceeds until it encounters a point $\tilde{x}$ satisfying $\|\hat{g}_{t}\|\leq 3\chi$, where $\chi$ is specified in (5). This point $\tilde{x}$ may lie near a saddle point with a large negative eigenvalue of the Hessian. To escape such a saddle point, the framework enters an escape procedure (steps 6–20), which performs $Q$ rounds of $\Gamma$-descent (steps 9–16).
@@ 5.3 Main Results for Gauss-PSGD Framework
[…]
@@
With probability at least $1-\omega/2$, for all $t\in[T]$, $\|\nu_{t}\|\leq C\sqrt{2\log\left(\frac{4T}{\omega}\right)}\psi\leq\chi$.
@@ 6.2 Error Rate Analysis for DP-SOSP with Ada-DP-SPIDER
    /* Using oracle $\mathcal{O}_{1}$ */
@@ 9    $\textsf{drift}\leftarrow 0$;
    /* Using oracle $\mathcal{O}_{2}$ */
@@ 7 Extension to Distributed SGD
The learning algorithm using Distributed Ada-DP-SPIDER can be viewed as an adaptive extension of the DIFF2 algorithm [36], which uses standard SPIDER and is limited to convergence to DP-FOSP under homogeneous data. To the best of our knowledge, our method is the first to achieve convergence to a DP-SOSP in a distributed setting with heterogeneous data.
@@ A.2 Privacy Preliminaries
Given any input data $D\in\mathcal{X}^{n}$ and a query function $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, the Gaussian mechanism $\mathcal{M}_{G}$ is defined as $q(D)+\nu$ where $\nu\sim\mathcal{N}(0,\sigma_{G}^{2}\mathbf{I}_{d})$. Let $\Delta_{2}(q)$ be the $\ell_{2}$-sensitivity of $q$, i.e., $\Delta_{2}(q)\coloneqq\sup_{D\sim D^{\prime}}\|q(D)-q(D^{\prime})\|_{2}$. For any $\sigma,\delta>0$, $\mathcal{M}_{G}$ guarantees $(\frac{\Delta_{2}(q)}{\sigma_{G}}\sqrt{2\log\frac{1.25}{\delta}},\delta)$-DP. That is, if we want the output of $q$ to be $(\epsilon,\delta)$-DP for any $0<\epsilon,\delta<1$, then $\sigma_{G}$ should be set to $\frac{\Delta_{2}(q)}{\epsilon}\sqrt{2\log\frac{1.25}{\delta}}$.
@@
Given target privacy parameters $0<\epsilon<1$ and $0<\delta<1$, to ensure $(\epsilon,\delta)$-DP over $k$-fold adaptive mechanisms, it suffices that each mechanism is $(\epsilon^{\prime},\delta^{\prime})$-DP, where $\epsilon^{\prime}=\frac{\epsilon}{2\sqrt{2k\ln(2/\delta)}}$ and $\delta^{\prime}=\frac{\delta}{2k}$.
@@ B.6 Proof of Lemma 5
Applying a union bound over $t\in[T]$ gives the desired result: with probability at least $1-\omega/2$, $\|\hat{g}_{t}-\nabla F(x_{t-1})\|\leq\chi$ for all $t$. ∎
@@ B.7 Proof of Lemma 6
By Lemma 5, with probability at least $1-\omega/2$, the gradient estimation error satisfies $\|\hat{g}_{t}-\nabla F(x_{t-1})\|\leq\chi$ for all $t\in[T]$. We analyze two cases based on whether the algorithm is in the escape phase.
@@
By Corollary 1, each $\Gamma$-descent succeeds with probability at least $1/8$, and we boost this to $1-\omega/2$ via the $Q$ independent repetitions in every escape procedure. By Lemma 2 with failure probability $\omega_{0}=\frac{\omega}{2N_{\text{saddle}}}$, we require:
@@ C.2 Proof of Lemma 8
By Lemma 5, with probability at least $1-\omega/2$, we have $\|\nabla F(x_{t-1})-\hat{g}_{t}\|\leq\chi$ for all $t$.
@@
$U\leq-\frac{|\mathcal{T}|\kappa}{6\eta}+5T\eta\chi^{2},$ (97)
@@ D.3 Proof of Theorem 4
Gradient Estimation Error. For any $\mathcal{S}_{j}$ and $x$, $\nabla\hat{f}_{\mathcal{S}_{j}}(x)-\nabla F_{j}(x)$ is zero-mean and follows $\operatorname{nSG}\left(\frac{2G}{\sqrt{n}}\right)$. By the $G$-Lipschitz assumption and norm-sub-Gaussian concentration (Lemma 11), we have with probability at least $1-\omega^{\prime}/8$:
@@
Also, since $\theta_{p}\sim\mathcal{N}(0,\sigma_{1}^{2}/m)$, standard Gaussian concentration (Lemma 10) gives, with probability at least $1-\omega^{\prime}/8$:
@@
Hessian Estimation Error. For any $j\in[m]$ and $z\in\mathcal{S}_{j}$, $\mathbb{E}[\nabla^{2}f(x_{p};z)-\nabla^{2}F_{j}(x_{p})]=0$, and $\|\nabla^{2}f(x_{p};z)-\nabla^{2}F_{j}(x_{p})\|_{2}\leq 2M$ (due to $M$-smoothness). That is, each empirical Hessian term is $2M$-bounded in operator norm. Applying the matrix Bernstein inequality (Lemma 14), and using the assumption $mn\geq\frac{4}{9}\log(8d/\omega^{\prime})$, we obtain with probability at least $1-\omega^{\prime}/8$:
@@
For the added noise, since $\mathbf{H}_{p}$ consists of symmetric Gaussian matrices with variance $\sigma_{2}^{2}/m$, Lemma 15 gives, with probability at least $1-\omega^{\prime}/8$:
@@
Verification for $x_{p}$. Combining the above estimates and using a union bound, with probability at least $1-\omega^{\prime}/2$, we have:
@@
$\geq-\left(\sqrt{\rho\alpha}+\text{(estimation error)}\right)$ (134)
@@
Hence, $x_{p}$ will be selected with probability at least $1-\omega^{\prime}/2$.